OpenStudy (anonymous):

Write a parameterization for the curve in the xy-plane. The line traced out from the point (11, −1) at t = 0 to the point (9, 8) at t = 1.

OpenStudy (anonymous):

Let x=a+mt y=b+nt be the parametric equations. Let t=0, and t=1 to find a,b,m,n. Voila!